OpenStudy (anonymous):

Find the unit rate.... $161/1 share ?

OpenStudy (anonymous):

anwser-$23/1 share

OpenStudy (sasogeek):

not sure what the question means... :/

OpenStudy (anonymous):

I have to fin the unit rate

OpenStudy (anonymous):

A unit rate is a ratio of two units.It expresses how many units of one quantity corresponds to one unit of another type of quantity.

jimthompson5910 (jim_thompson5910):

If you mean 161/7, then 161/7 = 23/1

OpenStudy (anonymous):

yea thats correct...thanks :) How did you know it was 7 and not 1 i made a mistake and you fixed it with out looking at my book....huh oh well thanks mister or miss

jimthompson5910 (jim_thompson5910):

well you gave the answer as 23/1 and I just made that connection

OpenStudy (anonymous):

How did you find out that 191/7 shares =23/1 ??????????????????????

jimthompson5910 (jim_thompson5910):

Divide both components of 161/7 by the GCF 7 So 161/7 = 23 and 7/7 = 1 This means that 161/7 = 23/1

OpenStudy (sasogeek):

23 x 7 =161 23=161/7 23/1=161/7